# Counting roots of unity on the graphs of Laurent series over non-Archimedean local fields

### v2

@@ 3.1. The ramified part
    If $|f|_{1}=1$ and $K(f,1)=k(f,1)$, then c_f = {2max(—K(f,1)—,1) if $\tilde{f}=0$ 2max(—K(f,1)—,— K(~f,1)—,— k(~f, 1)—,1) otherwise ,
@@
    If $|f_{k(f,1)}-b|<|\tilde{f}|_{1}$, then K(a,1)-k(a,1)=—k_1—(K(~f,1)-k(~f,1)).
@@
Let $K$ be a $p$-adic number field and $f\in\mathcal{A}_{K}[1,1]$. Assume that $f$ is not special. If $(\zeta,f(\zeta))$ is a torsion of $(K^{\times})^{2}$ of order $mp^{k}$ with $p\nmid m$, then $p^{k}$ is bounded by a constant only depending on $f$ and the ramification degree of $K$. More precisely,
@@
Fix a primitive $m$-th root of unity $\zeta_{m}$ and a primitive $e$-th root of unity $\zeta_{e}$. We define the auxiliary power series
$a(X):=\zeta_{m}^{(2)}\xi_{e}^{(2)}X^{k_{2}}-f(\zeta_{m}^{(1)}\xi_{e}^{(1)}X^{k_{1}})\in K(\zeta_{m},\zeta_{e})[[X,X^{-1}]].$
@@
Let $\eta_{m},\eta_{e}$ respectively denote a primitive $m$-th and a primitive $e$-th root of unity. Recall that by Lemma 2.11
@@
Galois conjugates of $\xi_{m}$ over $K(\eta_{m},\eta_{e})$. By Lemma 2.12, $\sigma a(x)=a(\sigma x)$ for all $\sigma\in\operatorname{Gal}(\bar{K}/K)$, so $a$ has at least that many zeros in $\operatorname{\mathbb{C}}_{p}$.
(iv) Let $N$ be the number of zeros of $a$ in $\operatorname{\mathbb{C}}_{p}$. Then, combining (ii) and (iii), we obtain

### v3

@@ 3.1. The ramified part
    If $|f|_{1}=1$ and $K(f,1)=k(f,1)$, then
    $c_{f}=\begin{cases}2\max(|K(f,1)|,1)&\text{ if $\tilde{f}=0$ }\\ 2\max(|K(f,1)|,|K(\tilde{f},1)|,|k(\tilde{f},1)|,1)&\text{ otherwise }\end{cases},$
@@
    If $|f_{k(f,1)}-b|<|\tilde{f}|_{1}$, then
    $K(a,1)-k(a,1)=|k_{1}|(K(\tilde{f},1)-k(\tilde{f},1)).$
@@
Let $K$ be a $p$-adic number field and $f\in\mathcal{A}_{K}[1,1]$. Assume that $f$ is not special. If $(\zeta,f(\zeta))$ is a torsion point of $(K^{\times})^{2}$ of order $mp^{k}$ with $p\nmid m$, then $p^{k}$ is bounded by a constant only depending on $f$ and the ramification degree of $K$. More precisely,
@@
Fix a primitive $m$-th root of unity $\eta_{m}$ and a primitive $e$-th root of unity $\eta_{e}$. We define the auxiliary power series
$a(X):=\zeta_{m}^{(2)}\xi_{e}^{(2)}X^{k_{2}}-f(\zeta_{m}^{(1)}\xi_{e}^{(1)}X^{k_{1}})\in K(\eta_{m},\eta_{e})[[X,X^{-1}]].$
@@
We will now bound the number of Galois conjucates of $\xi_{p^{k}}$ from below. Recall that by Lemma 2.11
@@
Galois conjugates of $\xi_{p^{k}}$ over $K(\eta_{m},\eta_{e})$. By Lemma 2.12, $\sigma a(x)=a(\sigma x)$ for all $\sigma\in\operatorname{Gal}(\bar{K}/K)$, so $a$ has at least that many zeros in $\operatorname{\mathbb{C}}_{p}$.
Let $N$ be the number of zeros of $a$ in $\operatorname{\mathbb{C}}_{p}$. Then, combining both bounds, we obtain